In eleven years, Lita will be 5 times as old as she was 9 years ago. How old is she?
Answer by nerdybill(7384)

In eleven years, Lita will be 5 times as old as she was 9 years ago. How old is she?
.
Let x = age of Lita
then
x+11 = Lita's age in eleven years
x-9 = Lita's age 9 years ago
.
x+11 = 5(x-9)
x+11 = 5x-45
11 = 4x-45
56 = 4x
14 years old = x